OpenStudy (skarlettrose186):

Eliza started her savings account with $100. Each month she deposits $25 into her account. Determine the average rate of change in Eliza's account from the 2nd month to the 10th month. Show all work for full credit.

OpenStudy (skarlettrose186):

I have this much m=months x= total amount x=100 + 25m Month Two: x = 100 + 25*2 x= 100 + 50 x= 150 Month Ten: x= 100 + 25*10 x= 100 + 250 x= 350

OpenStudy (wanell):

Relationship: x=100+25y where y is the amount of months After 2 months: x=100+25(2) x=150 After 10 months: x=100+25(10) x=350 Change between 2nd and 10th month is $200.

OpenStudy (wanell):

does this help

OpenStudy (skarlettrose186):

what would the equation for the rate of change be?